OpenStudy (twizttiez):

Variable y varies directly with variable x, and y = 16 when x = 12. Enter the constant of variation in lowest terms. ? y= ------x ? I think my answer is 4/3

OpenStudy (twizttiez):

@Data_LG2 this ? is more important

OpenStudy (anonymous):

you are correct (16/12) ---.(4/3)

OpenStudy (twizttiez):

y varies directly with variable x y = kx, where k is any constant y=16; x=12 16 = 12k k=16/12 = 4/3 (in lowest terms) I think

OpenStudy (anonymous):

\(\sf m=\large \frac{ \Delta y}{ \Delta x}=\frac{ 16}{ 12}=\frac{ 4}{ 3}\) yeah you're right (:

OpenStudy (twizttiez):

YAY! i have one more care to help?